The javascript on this page is used to preload the images for links and to swap the images used for links
go to content
layout image go to introduction
go to science datago to mission status reportgo to location / positiongo to 3d data mapsgo to moon VRMLgo to further reading
layout image
Layout imagego to home
go to projectgo to resultsGo to datavisgo to historyGo to scienceGo to educationGo to archivesgo to resources


Lunar Prospector Status Report #71

February 28, 1999 - 1:00 p.m. EST (10:00 a.m. PST)


The Lunar Prospector spacecraft is operating well and collecting good science data in the extended mission orbit. Four of five science instruments are on: the Gamma Ray Spectrometer (GRS), the Neutron Spectrometer (NS), the Magnetometer (MAG) and the Electron Reflectometer (ER). Data continues to be collected for the Doppler Gravity Experiment (DGE) as well.

Cycling of the spacecraft transmitter continues. The transmitter is powered off for occultations (when the spacecraft goes behind the moon and no data is received at Earth anyways) to reduce battery drain during propellant tank heater events. There was one instance (DOY 050) when the commands to turn off the transmitter were not sent becuase of station configuration problems, but there was no impact on the battery. On DOY 051, commands to turn the transmitter back on were delayed due to station configuration problems, and 28 mintues of data were lost.

On DOY 056, thrusters were fired to maintain the extended mission orbit for another 28 days. The timeline is given below (all times are GMT):

056/10:45 Turn A1,A2 thruster heaters on
056/10:50 Load maneuver parameter file
056/11:16 Fire part 1: A1/A2 for 43.5 sec; deltaV = 6.85 m/s
  Raise periapsis from 15.8 km to 40.5 km; used 0.55 kg propellant
056/11:24 Turn A1,A2 thruster heaters on
056/11:30 Load maneuver parameter file
056/11:55 Fire part 2: A1/A2 for 46.5 sec; deltaV = 7.34 m/s
  Set new periapsis - new orbit 15 x 45 km; used 0.59 kg propellant
056/12:04 Turn A1,A4 thruster heaters on
056/12:40 Load maneuver paramenter file
056/12:44 Fire reorientation maneuver: 7 0.2-sec pulses of A1/A4
  Precess spin axis 1.2 deg; target attitude 348/-89.1; 0.02 kg prop
056/12:48 Turn T1 thruster heater on
056/12:59 Load maneuver parameter file
056/13:02 Fire spin trim: T1 for 0.62 sec; delta w = 0.155 rpm
Target spin rate 12.00 rpm; used 0.003 kg propellant

On DOY 057, two commands were sent to reset the gain of the GRS instrument.

Preliminary results of the maneuver Thursday show excellent performance.

The next maneuver is scheduled for March 25.



Current spacecraft state (0000 GMT 2/26/99, DOY 057):

Orbit: 5048
Downlink: 3600 bps
Spin Rate: 12.00 rpm

Spin Axis Attitude (ecliptic):
Latitude: -89.2 deg
Longitude: 337.5 deg

Trajectory:

Periapsis Alt: 15.9 km
Apoapsis Alt: 44.1 km
Period: 111 min


Occultations: none
Eclipses: 47 minutes
Propellant Remaining: 14.78 kg